OpenStudy (anonymous):

A floor plan is given for the first floor of a new house. One inch represents 6 feet. If the walk-in closet is 1/4 in. by 1/3 in. on the floor plan, what are the actual dimensions?

OpenStudy (anonymous):

@mathstudent55 can you help with my last problem please?

OpenStudy (mandre):

6 ft is 6 * 12 = 72 inches so the scale is 1 inch : 72 inches. So multipy each by 72.

OpenStudy (anonymous):

I got answers that aren't in my choices

OpenStudy (mathstudent55):

You can also do it with both units. If 1 inch represents 6 feet, then every measurement you see in inches, multiply by 6 and the result is in feet.

OpenStudy (mathstudent55):

6 * 1/4 = ? 6 * 1/3 = ?

OpenStudy (anonymous):

i got it thanks!!

OpenStudy (mathstudent55):

If, on the other hand you do it by multiplying by 72, then your measurments will be in inches, so you need to divide them by 12 to convert to feet.

OpenStudy (mathstudent55):

Using the scale directly (1 inch : 6 ft): 6 * 1/4 = 1.5 ft 6 * 1/3 = 2 ft Using the inch scale (1 inch : 72 inches): 72 * 1/4 = 18 inches; 18/12 = 1.5 ft 72 * 1/3 = 24 inches; 24/12 = 2 ft. Either way you get 1.5 ft by 2 ft.